OpenStudy (anonymous):

Solve: -2x - 15 = 6x + 9 3 -3 -1 6

OpenStudy (igreen):

-2x - 15 = 6x + 9 Add 15 to both sides: -2x = 6x + 24 Subtract 6x to both sides: -8x = 24 Divide -8 to both sides: x = -3

OpenStudy (anonymous):

Solve: -5x = 30 6 -6 25 35

OpenStudy (texaschic101):

you can even check your answer to make sure -2x - 15 = 6x + 9 -2(-3) - 15 = 6(-3) + 9 6 - 15 = -18 + 9 -9 = -9 (correct) ============= -5x = 30 --- divide both sides by -5 x = 30/-5 x = -6

OpenStudy (anonymous):

Given the equation 3x - 3 = 6x + 1, select the reasoning that correctly solves for x. Add 3, subtract 3x, then divide by -3. Add 3, subtract 6x, then divide by 3. Subtract 1, subtract 3x, then divide by 3. Subtract 1, subtract 6x, then divide by -3. is itb

OpenStudy (anonymous):

@tejasvir

OpenStudy (anonymous):

@TexasBoy123

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

olve: -4(2x - 1) = -6(x + 2) - 2 -5 5 9 -9

OpenStudy (texaschic101):

3x - 3 = 6x + 1 --- subtract 1 3x - 4 = 6x -- subtract 3x -4 = 3x -- divide by 3 x = -4/3 =============== -4 (2x - 1) = -6 (x + 2) - 2 --- distribute through the parenthesis -8x + 4 = -6x - 12 - 2 --- combine like terms -8x + 4 = -6x - 14 -- subtract 4 -8x = -6x - 18 -- add 6x -2x = -18 -- divide by -2 x = -18/-2 x = 9